OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation.

OpenStudy (anonymous):

If you tell me the system I'm an expert at this.

OpenStudy (anonymous):

6x – y = 1 4x – 3y = –11

OpenStudy (anonymous):

Substitution method is super easy. :) Since it said find the equivalent to y in the first equation... you need to isolate y! -y = -6x + 1 y = 6x - 1 Now that you have your equation, just substitute that into the second equation. 4x -3(6x - 1) = -11 4x - 18x + 3 = -11 -14x + 3 = -11 -14x = -14 now divide x = 1 YOUR answer however would be y = 6x - 1